Write a quadratic equation in the variable x having x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0 solution: 3 only solution
Answer by stanbon(75887)

Write a quadratic equation in the variable x having x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0 solution: 3 only solution
--------
y = (x-3)(x-3)
---
y = x^2-6x+9
=========================
Cheers,
Stan H.